# Near-optimal Linear Predictive Clustering in Non-separable Spaces via Mixed Integer Programming and Quadratic Pseudo-Boolean Reductions

### v1

@@ 3.1 Preliminary Definitions and MIP formulation with Global Optimality
Since $\mathbf{Z}_{k,i,i}=1$ only if sample $i$ is assigned to cluster $k$, the residual $\mathbf{y}_{i}-\mathbf{X}_{i}\mathbf{w}_{k}$ is nonzero only for assigned samples. Consequently, the SSE for $\mathbf{w}_{k}$ is computed exclusively over samples in cluster $k$. LPC aims to determine the optimal $\mathbf{Z}_{k}$ and $\mathbf{w}_{k}$ such that the total SSE across all clusters is minimized.
@@ 3.2 Optimization of Regression Coefficients $\mathbf{w}$
Eliminating $\mathbf{Z}_{k}$ from $\mathbf{A}_{k}$ enables precomputing $\mathbf{A}^{\ast}$, reducing the optimization of $\mathbf{w}$ to a simple matrix multiplication, leading to a simplified MIP formulation. Specifically, substituting the $\mathbf{w}_{k}^{\ast}$ approximation back into the objective function in (1), we derive a quadratic form solely in terms of the cluster assignment variables $\mathbf{Z}_{k}$:
@@
Here, the only variable requiring optimization is the binary indicator $\mathbf{Z}_{k}$, while the regression coefficients $\mathbf{w}$ are implicitly optimized as given in (7). This transformation successfully reduces the MIQP formulation of LPC into a higher-order Pseudo-Boolean Optimization (PBO) problem (cf. App. B), which we reduce to a Quadratic PBO (QPBO) form in Sec. 3.4.
@@ 4.2 Experimental Setup
GlobalOpt follows the formulation in (2). Greedy optimization follows the algorithm proposed by chembu2023generalized. Since Greedy is sensitive to initialization, each experimental trial evaluates Greedy Optimization using 20 different initializations with a maximum of 100 iterations and reports the mean performance. We used the industry-standard solver Gurobi (gurobi), with an optimality gap fixed at 5%. Although LPC-NS-QPBO is formulated as a QPBO problem, we observed that the MIP solver in Gurobi dedicated PBO solver SCIP (BolusaniEtal2024ZR) (cf. App. E for performance evaluations). We rely on Gurobi’s automatic optimization of $M$ and were not able to outperform it with a customized constraint.
@@ 5 Experimental Results
All code to reproduce these results is available online. [^1] https://github.com/D3Mlab/LPC-NS
@@
Figure 4: RQ2 Noise in Target Variables Difference from the GlobalOpt objective and performance in recovering $\mathbf{w}^{\text{gt}}$ and $\mathbf{Z}^{\text{gt}}$ as the level of Gaussian noise in the target increases. Scatter points show mean results, with the solid line indicating the 95% confidence interval.
@@
Figure 6: RQ2 Proportion of Outliers The performance of different methods as the proportion of outliers in target variables increases. (cf. $K=3$ in App. F due to page constraints).
@@ 6 Conclusion
With the aim to develop scalable and near-optimal LPC optimization methods, we leverage properties of non-separable feature spaces to derive novel MIP and QPBO reductions from an existing globally optimal MIP formulation. Specifically, LPC-NS-MIP offers a more compact and efficient formulation than GlobalOpt while maintaining near-optimal performance in non-separable settings. LPC-NS-QPBO further enhances scalability via a Quadratic Pseudo-Boolean (QPBO) reduction. Extensive experiments show that LPC-NS-MIP consistently outperforms Greedy in optimizing the objective and recovering hidden cluster assignments and regression coefficients across varying noise, dimensions, and outliers for $K=2$ and $K=3$, while LPC-NS-QPBO remains highly competitive for $K=2$.

### v2

@@ 3.1 Preliminary Definitions and MIP formulation with Global Optimality
Since $\mathbf{Z}_{k,i,i}=1$ only if sample $i$ is assigned to cluster $k$, the residual $\mathbf{y}_{i}-\mathbf{X}_{i}\mathbf{w}_{k}$ is nonzero only for assigned samples. Consequently, the SSE for $\mathbf{w}_{k}$ is computed exclusively over samples in cluster $k$. LPC aims to determine the optimal $\mathbf{Z}_{k}$ and $\mathbf{w}_{k}$ such that the total SSE across clusters is minimized.
@@ 3.2 Optimization of Regression Coefficients $\mathbf{w}$
Eliminating $\mathbf{Z}_{k}$ from $\mathbf{A}_{k}$ enables precomputing $\mathbf{A}^{\ast}$, reducing the optimization of $\mathbf{w}$ to a matrix multiplication, leading to a simplified MIP formulation. Specifically, substituting the $\mathbf{w}_{k}^{\ast}$ back into (1), we derive a quadratic form solely in terms of the cluster assignment variables $\mathbf{Z}_{k}$:
@@
Here, the only variable requiring optimization is the binary indicator $\mathbf{Z}_{k}$, while the regression coefficients $\mathbf{w}$ are implicitly optimized as given in (7). This transformation successfully reduces the MIQP formulation of LPC into a higher-order Pseudo-Boolean Optimization (PBO) problem, which we reduce to a Quadratic PBO (QPBO) form in Sec. 3.4.
@@ 4.2 Experimental Setup
GlobalOpt follows the formulation in (2). Greedy optimization follows the algorithm proposed by chembu2023generalized. Since Greedy is sensitive to initialization, each experimental trial evaluates Greedy Optimization using 20 different initializations with a maximum of 100 iterations and reports the mean performance. We used the industry-standard solver Gurobi (gurobi), with an optimality gap fixed at 5%. Although LPC-NS-QPBO is formulated as a QPBO problem, we observed that the MIP solver in Gurobi outperforms the PBO solver SCIP (BolusaniEtal2024ZR) (cf. App. E for performance evaluations). We rely on Gurobi’s automatic optimization of $M$ and were unable to outperform it using customized constraints.
@@ 5 Experimental Results
Figure 4: RQ2 Noise in Target Variables The performance difference across different methods with the 95% confidence interval as the level of Gaussian noise in the target increases.
@@
Figure 6: RQ2 Proportion of Outliers The performance of different methods as the proportion of outliers in target variables increases. ($K=3$ cf. App. F.1 due to page constraints).
@@ 6 Conclusion
With the aim to develop scalable and near-optimal LPC optimization methods, we leverage properties of non-separable feature spaces to derive novel MIP and QPBO reductions from an existing globally optimal MIP formulation. Specifically, LPC-NS-MIP offers a more compact and efficient formulation than GlobalOpt while maintaining near-optimal performance in non-separable settings. LPC-NS-QPBO further enhances scalability via a QPBO reduction. Extensive experiments show that LPC-NS-MIP consistently outperforms Greedy in optimizing the objective and recovering hidden cluster assignments and regression coefficients across varying noise, dimensions, and outliers, while LPC-NS-QPBO remains highly competitive for $K=2$.